OpenStudy (anonymous):

Auditions for a new, upcoming band in Chicago were held with 72 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 34, how many women auditioned?

OpenStudy (radar):

X = number of men Y = number of woman 2X-3Y=34 X+Y=72 multiply 2nd equation by 3 3X+3Y=216 add ti first equation: 2X-3Y=34 ------------ 5X=250 X=50 Y=22